# A nonparametric approach to practical identifiability of nonlinear mixed effects models

### v1

@@ Acknowledgments
Here, we set $\sigma^{2}=0.025$ and considered population sizes $n=5,50,100,$ and 200200200 to illustrate how the identifiability of the population estimates depends on the amount of data available for calibration.
@@
We compute the log-likelihood in Eq. (5) via Monte-Carlo integration to estimate the population-level parameters. In Figure 6, we plot the log-likelihood function evaluated at 800 uniformly sampled parameter pairs $(\mu_{a},\mu_{b})$ to estimate the population-level parameters of the simple exponential growth model in Eq. (4) for population sizes of $n=5,20,50,200$. As the model is clearly unidentifiable at the individual replicate level, it is unsurprising to see that the likelihood does not accurately capture the population-level parameters $\mu_{a}$ and $\mu_{b}$ for small population sizes $n=5$. However, as the population size increases, especially to $n=50$ and $n=200$, the likelihood begins to more accurately identify the underlying values of $\mu_{a}$ and $\mu_{b}$. Figure  6 thus illustrates the analytical results of Lavielle and Aarons, [40], and demonstrates that hierarchical parameter estimation can facilitate the identification of model parameters that would otherwise be unidentifiable.
@@
Figure 6: The log-likelihood distribution as a function of the mean values of the exponential distributions for the growth rates $a$ and $b$, $\mu_{a}$ and $\mu_{b}$. Panels A-D show the likelihood for $n=5,20,50$, and 200200200 virtual replicates. The true value of the mean growth rates, $\mu_{a}$ and $\mu_{b}$, are plotted as a diamonds. Since the inference problem is identical after relabelling $a$ and $b$, the pairs $(\mu_{a},\mu_{b})$ and $(\mu_{b},\mu_{a})$ both represent true values. The best 4000 samples from the likelihood are shown as filled circles. The likelihood has a reflection symmetry through the line $\mu_{a}=\mu_{b}$, shown as a grey dashed line, which corresponds to the relabelling of $a$ and $b$.

### v2

@@ Acknowledgments
Here, we set $\sigma^{2}=0.025$ and considered population sizes $n=5,50,100,$ and $200$ to illustrate how the identifiability of the population estimates depends on the amount of data available for calibration.
@@
We compute the log-likelihood in Eq. (5) via Monte-Carlo integration to estimate the population-level parameters. In Figure 7, we plot the log-likelihood function evaluated at 800 uniformly sampled parameter pairs $(\mu_{a},\mu_{b})$ to estimate the population-level parameters of the simple exponential growth model in Eq. (4) for population sizes of $n=5,20,50,200$. As the model is clearly unidentifiable at the individual replicate level, it is unsurprising to see that the likelihood does not accurately capture the population-level parameters $\mu_{a}$ and $\mu_{b}$ for small population sizes $n=5$. However, as the population size increases, especially to $n=50$ and $n=200$, the likelihood begins to more accurately identify the underlying values of $\mu_{a}$ and $\mu_{b}$. Figure  7 thus illustrates the analytical results of Lavielle and Aarons, [45], and demonstrates that hierarchical parameter estimation can facilitate the identification of model parameters that would otherwise be unidentifiable.
@@
Figure 7: The log-likelihood distribution as a function of the mean values of the exponential distributions for the growth rates $a$ and $b$, $\mu_{a}$ and $\mu_{b}$. Panels A-D show the likelihood for $n=5,20,50$, and $200$ virtual replicates. The true value of the mean growth rates, $\mu_{a}$ and $\mu_{b}$, are plotted as a diamonds. Since the inference problem is identical after relabelling $a$ and $b$, the pairs $(\mu_{a},\mu_{b})$ and $(\mu_{b},\mu_{a})$ both represent true values. The best 4000 samples from the likelihood are shown as filled circles. The likelihood has a reflection symmetry through the line $\mu_{a}=\mu_{b}$, shown as a grey dashed line, which corresponds to the relabelling of $a$ and $b$.
Further results on the Friberg Model
Here, we present additional results on the identifiability of the Friberg model. We recall that we identified three distinct local minima of the log-likelihood and concluded that fits 1-3 correspond to the global minima of the log-likelihood. In Figure 8, we compare the estimated population distributions for $N_{0},k_{tr},k_{prol}$ and $\log_{10}(EC_{50})$ against the known population distributions used to generate the synthetic data. In general, the estimated distributions for fits 1-3 recapture the known underlying distributions. Taken together with our analysis in the main text, we conclude that the Friberg model is identifiable at the population level and the estimated population distributions recapture the underlying true distributions.
Figure 8: Comparison between the estimated and known parameter distributions from the the Friberg model. Panel A compares the 3 parameter estimates in cluster 1 against the known parameter distributions used to generate the synthetic data for each of the four model parameters. Panel B shows the overlapping index between each of the pairwise comparisons of the 3 parameter estimates in cluster 1 against the known parameter distributions. Here, for each of the $k=1,2,3,4$ model parameters, pairs of fits $(i,j)$ with an overlapping index $o_{i,j,k}>0.5$ are represented by the shades of violet, while pairs with $o_{i,j,k}<0.5$ are represented by shades of orange.
In the previous analysis, we considered $N=15$ participants. To analyse if the Friberg model remains practically identifiable at the population level for a smaller number of participants, we consider $N=5$ virtual participants. We once again considered 100 initial estimates for the model parameters and fit the synthetic data from these initial estimates. As in the main text, we consider the 10 best fits to the data, as measured by AIC (or log-likelihood). As in the main text, the model captured the individual dynamics for all 5 virtual participants. We therefore considered the individual and population estimates for the model parameters. These parameter estimates are shown in Figure 9. Broadly speaking, the Kolmogorov-Smirnov test does not identify significant differences between the distributions of individual estimates for $N_{0},\ k_{tr}$ and $k_{prol}$. However, there are some pairwise significant differences between the distributions of individual estimates for $\log_{10}(EC_{50})$. These results are mirrored in the population distributions, where there are large overlaps between $N_{0},\ k_{tr}$ and $k_{prol}$. Conversely, some pairs of population distributions of $\log_{10}(EC_{50})$ do not have large overlap. Consequently, these results suggest that the population distribution of $\log_{10}(EC_{50})$ is poorly identified when fitting the model to $N=5$ participants.
Figure 9: Assessing practical identifiability of the Friberg model for $N=5$ virtual participants. Panel A shows the distribution of the 15 individual estimates for the model parameters $N_{0},k_{tr},k_{prol}$ and $\log(EC_{50})$ for each of the 10 best population fits as dots within each violin plot. Panel B shows the $p$-value corresponding to the pairwise Kolmogorov-Smirnov test between all 10 best fits for each of the four model parameters. Here, $p$-values below the significance threshold of $0.05$ are plotted as blue squares. Panel C shows the 10 estimated population distributions for each of the four model parameters. Panel D shows the overlapping index between each of the pairwise comparisons of the 10 best fits. Here, for each of the $k=1,2,3,4$ model parameters, pairs of fits $(i,j)$ with an overlapping index $o_{i,j,k}>0.5$ are represented by the shades of violet, while pairs with $o_{i,j,k}<0.5$ are represented by shades of orange and the three clusters are again highlighted by white boxes.